# Recurrence Relations and Dispersive Techniques for Precision Multi-Loop Calculations

### v1

@@ I Introduction
During the following decade, Refs. (Hollik1990, ; Denner1993, ) produced comprehensive reviews and explicit one-loop calculations relevant to LEP precision physics, codifying gauge-invariant renormalization schemes and parameter definitions. These works became standard references for both theoretical and computational approaches to radiative corrections. The extension to two-loop order demanded a deeper algebraic understanding of Feynman integrals. Refs. (Tarasov96, ; Tarasov97, ) introduced dimension-recurrence and propagator-power-reduction identities that relate integrals in $d$ and $d\pm 2$ dimensions. These relations enable systematic reduction of higher-rank and higher-dimensional integrals to a minimal set of master integrals. This formalism remains foundational for modern two-loop reduction algorithms and underpins most symbolic-manipulation packages used today. Concurrently, in (PLB91, ) a complementary algebraic framework was developed for reducing tensor Feynman integrals to scalar ones using dimension shifting recurrence relations.
@@
Analytic evaluation of increasingly complex diagrams soon became infeasible due to the proliferation of mass scales, external invariants, and threshold singularities. This challenge prompted the development of semi-analytical and numerical approaches such as sector-decomposition methods, differential-equation (DE) systems for master integrals, and dispersion-relation techniques (Lee, ; SmrBook, ; Kotikov, ). A major step was the differential-equation approach for master integrals, in which systems of linear equations in kinematic invariants are solved either analytically in canonical ($\varepsilon$-form) basis or numerically using series expansions. In (Wasser, ; Henn, ) this framework was refined by introducing canonical-basis and uniform-weight formulations, improving both analytic transparency and numerical stability.
@@
In (AA-1, ; AA-2, ; AA-3, ; AA-4, ; AA-5, ), it was developed a comprehensive two-loop framework, specifically for polarized Møller scattering: an essential channel for future parity-violation experiments. The group systematically advanced from reducible two-loop and quadratic one-loop contributions to two-loop irreducible self-energies, vertex and box calculations. These results quantified higher-order electroweak effects in parity-violating asymmetries, establishing reliable theoretical uncertainties at the sub-percent level-critical for the forthcoming MOLLER experiment.
@@
Beyond differential-equation and sector-decomposition strategies, a distinct line of development has been the dispersive approach developed in (AA1, )-(AB19a, ). This methodology expresses multi-point Passarino-Veltman functions in a two-point-function basis, replacing sub-loop insertions with effective propagators represented by dispersion integrals. The result is a semi-analytical bridge between purely analytic amplitude reductions and purely numerical integration techniques. The dispersive framework preserves key analytic properties — threshold behavior, unitarity cuts, and gauge invariance — while reducing computational demands. It is especially attractive for low-energy observables where delicate cancellations between diagrams require high numerical precision. Furthermore, real-experiment implementation often involves acceptance and energy-threshold cuts that significantly affect radiative corrections; the dispersive formulation allows these to be incorporated naturally at the numerical-integration stage.
@@
Parallel developments in the phenomenological sector culminated in partial two-loop electroweak predictions for key observables. In Refs. (Dubovyk1, ) and (Freitas1a, ) produced the full set of fermionic and bosonic two-loop corrections to Z-boson observables, while (Erler, ) addressed hadronic effects in Møller scattering at NNLO. More recently, (Schw, ) achieved the analytic evaluation of electroweak double-box integrals relevant for Møller processes.
@@
These advances collectively establish a robust infrastructure for high-precision electroweak phenomenology. A comprehensive review on updated measurements and higher-order theoretical corrections is available at (Erler1, ) and (PDG2024, ). Despite impressive progress, complete NNLO electroweak results remain available only for selected processes due to the technical difficulty of two-loop calculations. Closed-fermion-loop NNLO corrections have been achieved for several key observables, but the general problem of fully automated two-loop amplitude generation, reduction, and evaluation remains open. Semi-numerical strategies, such as the dispersive and differential-equation methods, have proven especially effective, balancing analytical control with numerical tractability. The growing complexity of precision calculations arises from the need to handle diagrams with both more external legs and additional loop orders. Each loop introduces new mass and momentum scales, overlapping ultraviolet and infrared divergences, and complicated threshold structures, dramatically increasing algebraic and numerical challenges. At one loop, PV-style analytic reductions allow tensor integrals to be expressed through a small set of scalar functions, often in closed form. At two loops and beyond, however, the explosion of topologies and kinematic invariants typically renders complete analytic evaluation impractical. To address these challenges, the dispersive method reformulates a multi-loop integral as a sequence of nested one-loop integrals via spectral representations. This effectively transforms a two-loop problem into integrals over well-behaved spectral densities, isolating singular behavior in analytically controlled functions that can be integrated numerically with high stability. By combining dimensional shifting, recurrence relations, and dispersion theory, the approach achieves algebraic reduction to a minimal set of independent integrals while maintaining analytic transparency.
@@ II.3 The dispersion approach
${\rm Im}\left[{\rm i}^{-1}J^{(2)}(2-2\varepsilon;1,1)\right]_{s}=-2\pi^{1-\varepsilon}\frac{\Gamma(1-\varepsilon)}{\Gamma(1-2\varepsilon)}\;\frac{\pi}{\sqrt{-\Delta_{s}}}\left(\frac{s}{-\Delta_{s}}\right)^{\varepsilon}\;$ (42)
@@
$\frac{\partial}{\partial s}\left\{s^{-\varepsilon}{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2-2\varepsilon;1,1\right)\right]_{s}\right\}=-2\pi^{1-\varepsilon}\frac{\Gamma(1-\varepsilon)}{\Gamma(1-2\varepsilon)}\;\pi\;\left(\frac{\partial\Delta_{s}}{\partial s}\right)\left(\frac{\partial}{\partial\Delta_{s}}(-\Delta_{s})^{-1/2-\varepsilon}\right)\;,$
$\frac{\partial}{\partial m_{2}^{2}}{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2-2\varepsilon;1,1\right)\right]_{s}=-2\pi^{1-\varepsilon}\frac{\Gamma(1-\varepsilon)}{\Gamma(1-2\varepsilon)}\;\pi\;s^{\varepsilon}\left(\frac{\partial\Delta_{s}}{\partial m_{2}^{2}}\right)\left(\frac{\partial}{\partial\Delta_{s}}(-\Delta_{s})^{-1/2-\varepsilon}\right)\;.$
@@ III Numerical Examples
Figure 4: Numerical results for the four-point function $D_{0}$ ($k_{2}^{2}=-1.5\penalty 10000\ {\rm GeV}^{2}$, $k_{3}^{2}=-2.5\penalty 10000\ {\rm GeV}^{2}$, $k_{4}^{2}=m_{4}^{2}\penalty 10000\ {\rm GeV}^{2}$, $(k_{1}\cdot k_{3})=4.0\penalty 10000\ {\rm GeV}^{2}$, $(k_{2}\cdot k_{3})=-1.0\penalty 10000\ {\rm GeV}^{2}$ $m_{1}=1.5\penalty 10000\ {\rm GeV}$, $m_{2}=0.5\penalty 10000\ {\rm GeV}$, $m_{3}=2.0\penalty 10000\ {\rm GeV}$, $m_{4}=2.5\penalty 10000\ {\rm GeV}$). The functions $D_{2l,n_{1},n_{2},n_{3}}$ are defined in Eq. (15). Crossed dots are results based on this work and solid line is produced from Collier library.
@@ IV Roadmap to two-Loop calculations
As we can see from the previous chapters we have successfully represented one-loop (up to multiplicity four) integrals with an arbitrary tensor rank using recurrence and dispersive methods. In addition we where able to reduce higher multiplicity PV functions to two-point result. Finally, we have adopted dispersive technique introduced in (AA1, )-(AB19, ) to subtracted two-point functions. Now we have analytical results for PV functions with polynomial terms in $k^{2}$ and dispersive term carrying propagator like structure $\propto\frac{1}{(s-k^{2}-i0)}$. This particular representation is most valuable for applications in two-loop calculations for any possible particle physics models. First, if we consider $\left(j+2\right)$-point Feynman graph as an insertion (index $\left(j+2\right)$ means that we have $j$ number of external and two internal legs in the insertion) into the two-loop topology (see Fig. 5), then polynomial terms in external or second loop momenta will be a part of the numerator algebra and dispersive contribution will be treated as an additional propagator in the second loop integral.
Figure 5: The two-loop NN-point diagram ($k_{1:j}\equiv k_{1}+\ldots+k_{j}$).
@@
and first perform integration over the loop momentum $q_{2}$. That results in $J^{(3)}\left(D;1,1,1\right)$:
$J^{(3)}\left(D;1,1,1\right)=$ $\int\frac{{\rm d}^{D}q_{2}}{\left[q_{2}^{2}-m_{5}^{2}\right]\left[\left(q_{1}+q_{2}\right)^{2}-m_{1}^{2}\right]\left[\left(q_{1}+q_{2}-k\right)^{2}-m_{2}^{2}\right]}.$ (65)
Using Eq. (II.1), we get
@@
which effectively gives us a reduction of the three-point integral to a two-point one. Applying recursive approach outlined in Chapter II, we can lower second power of the last propagator in Eq.(69) and arrive to the subtracted $\bar{J}^{(2)}\left(2-2\varepsilon;1,1\right)$ plus terms containing $J^{(2)}\left(2-2\varepsilon;1,0\right)$ and $J^{(2)}\left(2-2\varepsilon;0,1\right)$ (see Eq. (32)). For the subtracted $\bar{J}^{(2)}\left(2-2\varepsilon;1,1\right)$, we can apply dispersive representation stemming from Eqs. (41)–(42). Finally, for $D=4-2\varepsilon$ and $\varepsilon\to 0$ Eq. (65) can be written in the following form:
$J^{(3)}\left(4-2\varepsilon;1,1,1\right)$ $=\frac{1}{2}\int\limits_{0}^{1}{\rm d}x\left\{\frac{\rm{i}\pi^{2-\epsilon}}{m_{5}^{2}-m_{12x}^{2}}\ln\frac{m_{12x}^{2}}{m_{5}^{2}}+\left[\left(q_{1}-\bar{x}k\right)^{2}+m_{5}^{2}-m_{12x}^{2}\right]\pi\bar{J}_{1}^{(2)}\left(2-2\varepsilon;1,1\right)\right\},$ (70)
where the function $\bar{J}_{1}^{(2)}\left(2-2\varepsilon;1,1\right)$ is defined in Eq. (41):
$\bar{J}_{1}^{(2)}\left(2-2\varepsilon;1,1\right)=\frac{{\rm i}}{\pi}\;\int\limits_{(m_{5}+m_{12x})^{2}}^{\infty}\frac{{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2-2\varepsilon;1,1\right)\right]_{s}}{s\left[\left(q_{1}-\bar{x}k\right)^{2}-s+{\rm i}0\right]}\;.$
Here, $m_{12x}^{2}=xm_{1}^{2}+\bar{x}m_{2}^{2}-x\bar{x}k^{2}$ and the imaginary part of ${\rm i}^{-1}J^{(2)}(2-2\varepsilon;1,1)$ at $\varepsilon=0$ has a simple structure:
$\rm{Im}\left[{\rm i}^{-1}J^{(2)}\left(2;1,1\right)\right]_{s}=-\frac{2\pi^{2}}{\sqrt{\left(s-m_{12x}^{2}-m_{5}^{2}\right)^{2}-4m_{12x}^{2}m_{5}^{2}}},$
In Eq. (70) we did not retain linear in $\varepsilon$ terms, since the insertion $J^{(3)}\left(D;1,1,1\right)$, and the entire two-loop integral are UV-finite. At this point, we are ready to complete integration over second loop momentum $q_{1}$:
@@
$-\frac{1}{\pi}\int\limits_{(m_{5}+m_{12x})^{2}}^{\infty}{\rm d}s\;\frac{{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2;1,1\right)\right]_{s}}{s}\left[\left(s+m_{5}^{2}-m_{12x}^{2}\right)C_{0}+B_{0}\right]\Biggr\},$ (71)
@@
$U^{(2)}=-\frac{1}{2\pi}\int\limits_{0}^{1}{\rm d}x\int\limits_{(m_{5}+m_{12x})^{2}}^{\infty}\frac{{\rm d}s}{s}\left(s+m_{5}^{2}-m_{12x}^{2}\right)\;{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2;1,1\right)\right]_{s}C_{0}.$ (72)

### v2

@@ I Introduction
During the following decade, Refs. (Hollik1990, ; Denner1993, ) produced comprehensive reviews and explicit one-loop calculations relevant to CERN’s Large Electron-Positron (LEP) collider precision physics, codifying gauge-invariant renormalization schemes and parameter definitions. These works became standard references for both theoretical and computational approaches to radiative corrections. The extension to two-loop order demanded a deeper algebraic understanding of Feynman integrals. Refs. (Tarasov96, ; Tarasov97, ) introduced dimension-recurrence and propagator-power-reduction identities that relate integrals in $d$ and $d\pm 2$ dimensions. These relations enable systematic reduction of higher-rank and higher-dimensional integrals to a minimal set of master integrals. This formalism remains foundational for modern two-loop reduction algorithms and underpins most symbolic-manipulation packages used today. Concurrently, in (PLB91, ) a complementary algebraic framework was developed for reducing tensor Feynman integrals to scalar ones using dimension shifting recurrence relations (see also (GG, ; BGHPS, ; EGZ, ) and references therein).
@@
Analytic evaluation of increasingly complex diagrams soon became infeasible due to the proliferation of mass scales, external invariants, and threshold singularities. This challenge prompted the development of semi-analytical and numerical approaches such as sector-decomposition methods, differential-equation (DE) systems for master integrals, and dispersion-relation techniques (Lee, ; SmrBook, ; Kotikov, ; BSch, ; BSSch, ). A major step was the differential-equation approach for master integrals, in which systems of linear equations in kinematic invariants are solved either analytically in canonical ($\varepsilon$-form) basis or numerically using series expansions. In (Wasser, ; Henn, ) this framework was refined by introducing canonical-basis and uniform-weight formulations, improving both analytic transparency and numerical stability.
@@
A comprehensive two-loop framework was developed in (AA-1, ; AA-2, ; AA-3, ; AA-4, ; AA-5, ), specifically for polarized Møller scattering: an essential channel for future parity-violation experiments. The group systematically advanced from reducible two-loop and quadratic one-loop contributions to two-loop irreducible self-energies, vertex and box calculations. These results quantified higher-order electroweak effects in parity-violating asymmetries, establishing reliable theoretical uncertainties at the sub-percent level-critical for the forthcoming MOLLER experiment.
@@
Beyond differential-equation and sector-decomposition strategies, a distinct line of development has been the dispersive approach introduced in (AA1, )-(AB19a, ). This methodology expresses multi-point Passarino-Veltman functions in a two-point-function basis, replacing sub-loop insertions with effective propagators represented by dispersion integrals. The result is a semi-analytical bridge between purely analytic amplitude reductions and purely numerical integration techniques. The dispersive framework preserves key analytic properties — threshold behavior, unitarity cuts, and gauge invariance — while reducing computational demands. It is especially attractive for low-energy observables where delicate cancellations between diagrams require high numerical precision. Furthermore, real-experiment implementation often involves acceptance and energy-threshold cuts that significantly affect radiative corrections; the dispersive formulation allows these to be incorporated naturally at the numerical-integration stage. An example of an application of the dispersive approach to multi-loop “banana" integrals can be found in Ref. ChYCh . Another example is the precision studies of the Higgs boson at future colliders, which requires the inclusion of NNLO electroweak corrections. In (FR1a, ), the planar and non-planar double-box topologies with multiple massive propagators in the loops are computed numerically by transforming one of the sub-loops using Feynman parametrization and a dispersion representation. This was followed by (FR1b, ), which provides a complete calculation of the NNLO electroweak corrections involving closed fermion loops.
@@
Parallel developments in the phenomenological sector culminated in partial two-loop electroweak predictions for key observables. In Refs. (Dubovyk1, ) and (Freitas1a, ) it was produced the full set of fermionic and bosonic two-loop corrections to Z-boson observables, while (Erler, ) addressed hadronic effects in Møller scattering at NNLO. More recently, (Schw, ) achieved the analytic evaluation of electroweak double-box integrals relevant for Møller processes.
@@
These advances collectively establish a robust infrastructure for high-precision electroweak phenomenology. The comprehensive review on updated measurements and higher-order theoretical corrections are available at (Erler1, ) and (PDG2024, ). Despite impressive progress, complete NNLO electroweak results remain available only for selected processes due to the technical difficulty of two-loop calculations. Closed-fermion-loop NNLO corrections have been achieved for several key observables, but the general problem of fully automated two-loop amplitude generation, reduction, and evaluation remains open. Semi-numerical strategies, such as the dispersive and differential-equation methods, have proven especially effective, balancing analytical control with numerical tractability. The growing complexity of precision calculations arises from the need to handle diagrams with both more external legs and additional loop orders. Each loop introduces new mass and momentum scales, overlapping ultraviolet and infrared divergences, and complicated threshold structures, dramatically increasing algebraic and numerical challenges. At one loop, PV-style analytic reductions allow tensor integrals to be expressed through a small set of scalar functions, often in closed form. At two loops and beyond, however, the explosion of topologies and kinematic invariants typically renders complete analytic evaluation impractical. To address these challenges, the dispersive method reformulates a multi-loop integral as a sequence of nested one-loop integrals via spectral representations. This effectively transforms a two-loop problem into integrals over well-behaved spectral densities, isolating singular behavior in analytically controlled functions that can be integrated numerically with high stability. By combining dimensional shifting, recurrence relations, and dispersion theory, the approach achieves algebraic reduction to a minimal set of independent integrals while maintaining analytic transparency.
@@ II.3 The dispersion approach
${\rm Im}\left[{\rm i}^{-1}J^{(2)}(2-2\varepsilon;1,1)\right]_{s}=2\pi^{1-\varepsilon}\frac{\Gamma(1-\varepsilon)}{\Gamma(1-2\varepsilon)}\;\frac{\pi}{\sqrt{-\Delta_{s}}}\left(\frac{s}{-\Delta_{s}}\right)^{\varepsilon}\;$ (42)
@@
$\frac{\partial}{\partial s}\left\{s^{-\varepsilon}{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2-2\varepsilon;1,1\right)\right]_{s}\right\}=2\pi^{1-\varepsilon}\frac{\Gamma(1-\varepsilon)}{\Gamma(1-2\varepsilon)}\;\pi\;\left(\frac{\partial\Delta_{s}}{\partial s}\right)\left(\frac{\partial}{\partial\Delta_{s}}(-\Delta_{s})^{-1/2-\varepsilon}\right)\;,$
$\frac{\partial}{\partial m_{2}^{2}}{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2-2\varepsilon;1,1\right)\right]_{s}=2\pi^{1-\varepsilon}\frac{\Gamma(1-\varepsilon)}{\Gamma(1-2\varepsilon)}\;\pi\;s^{\varepsilon}\left(\frac{\partial\Delta_{s}}{\partial m_{2}^{2}}\right)\left(\frac{\partial}{\partial\Delta_{s}}(-\Delta_{s})^{-1/2-\varepsilon}\right)\;.$
@@ III Numerical Examples
Figure 4: Numerical results for the four-point function $D_{0}$ ($k_{2}^{2}=-1.5\penalty 10000\ {\rm GeV}^{2}$, $k_{3}^{2}=-2.5\penalty 10000\ {\rm GeV}^{2}$, $k_{4}^{2}=m_{4}^{2}$, $(k_{1}\cdot k_{3})=4.0\penalty 10000\ {\rm GeV}^{2}$, $(k_{2}\cdot k_{3})=-1.0\penalty 10000\ {\rm GeV}^{2}$ $m_{1}=1.5\penalty 10000\ {\rm GeV}$, $m_{2}=0.5\penalty 10000\ {\rm GeV}$, $m_{3}=2.0\penalty 10000\ {\rm GeV}$, $m_{4}=2.5\penalty 10000\ {\rm GeV}$). The functions $D_{2l,n_{1},n_{2},n_{3}}$ are defined in Eq. (15). Crossed dots are results based on this work and solid line is produced from Collier library.
@@ IV Roadmap to two-Loop calculations
As we can see from the previous chapters we have successfully represented one-loop (up to multiplicity four) integrals with an arbitrary tensor rank using recurrence and dispersive methods. In addition we where able to reduce higher multiplicity PV functions to two-point result. Finally, we have adopted dispersive technique introduced in (AA1, )-(AB19, ) to subtracted two-point functions. Now we have analytical results for PV functions with polynomial terms in $k^{2}$ and dispersive term carrying propagator like structure $\propto\frac{1}{(s-k^{2}-i0)}$. This particular representation is most valuable for applications in two-loop calculations for any possible particle physics models. It is worth mentioning that in this chapter we will consider only planar topology for the two-loop graphs. In the case of non-planar graphs, we can re-arrange momenta in the loop integral and reduce result to the dispersive representation considered in (AACR, ). First, if we consider $\left(j+2\right)$-point Feynman graph as an insertion (index $\left(j+2\right)$ means that we have $j$ number of external and two internal legs in the insertion) into the two-loop planar topology (see Fig. 5), then polynomial terms in external or second loop momenta will be a part of the numerator algebra and dispersive contribution will be treated as an additional propagator in the second loop integral.
Figure 5: The two-loop planar NN-point diagram ($k_{1:j}\equiv k_{1}+\ldots+k_{j}$).
@@
and first perform integration over the loop momentum $q_{2}$. That results in a three-point integral
$J^{(3)}\left(D;1,1,1\right)\Bigr|_{\footnotesize{\begin{array}[]{l}p_{1}=0,p_{2}=q_{1},p_{3}=q_{1}-k\\[-8.53581pt] m_{1}^{2}\leftrightarrow m_{5}^{2},m_{2}^{2}\leftrightarrow m_{1}^{2},m_{3}^{2}\leftrightarrow m_{2}^{2}\end{array}}}$ $\!\!\!\!=\int\frac{{\rm d}^{D}q_{2}}{\left[q_{2}^{2}\!-\!m_{5}^{2}\right]\left[\left(q_{1}+q_{2}\right)^{2}\!-\!m_{1}^{2}\right]\left[\left(q_{1}+q_{2}-k\right)^{2}\!-\!m_{2}^{2}\right]}.$ (67)
Using Eq. (II.1) (with $\bar{x}=1-x$), we get
@@
which effectively gives us a reduction of the three-point integral to a two-point one. Here it is assumed that $D=4-2\varepsilon$. Applying recursive approach outlined in Chapter II, we can lower second power of the last propagator in Eq.(71) and arrive to the subtracted $\bar{J}_{1}^{(2)}\left(2-2\varepsilon;1,1\right)$ plus tadpole terms containing $J^{(2)}\left(2-2\varepsilon;1,0\right)$ and $J^{(2)}\left(2-2\varepsilon;0,1\right)$ (see Eq. (32)). In this way, Eq. (67) can be written in the following form:
$J^{(3)}\left(4-2\varepsilon;1,1,1\right)=\frac{1}{2}\int\limits_{0}^{1}{\rm d}x\Biggl\{$ ${\rm i}\pi^{2-\varepsilon}\Gamma(\varepsilon)\;\frac{(m_{5}^{2})^{-\varepsilon}-(m_{12x}^{2})^{-\varepsilon}}{m_{5}^{2}-m_{12x}^{2}}$
$-\left[\left(q_{1}-\bar{x}k\right)^{2}+m_{5}^{2}-m_{12x}^{2}\right]\pi\bar{J}_{1}^{(2)}\left(2-2\varepsilon;1,1\right)\Biggr\},$ (72)
with $m_{12x}^{2}=xm_{1}^{2}+\bar{x}m_{2}^{2}-x\bar{x}k^{2}$. For the subtracted $\bar{J}_{1}^{(2)}\left(2-2\varepsilon;1,1\right)$, we can apply dispersive representation stemming from Eqs. (41)–(42),
$\bar{J}_{1}^{(2)}\left(2-2\varepsilon;1,1\right)=\frac{{\rm i}}{\pi}\;\int\limits_{(m_{5}+m_{12x})^{2}}^{\infty}{\rm d}s\;\frac{{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2-2\varepsilon;1,1\right)\right]_{s}}{s\left[s-\left(q_{1}-\bar{x}k\right)^{2}-{\rm i}0\right]}\;.$
Since the insertion $J^{(3)}\left(D;1,1,1\right)$, and the entire two-loop integral are UV-finite, we can consider the limit $\varepsilon\to 0$:
$J^{(3)}\left(4;1,1,1\right)$ $=\frac{1}{2}\int\limits_{0}^{1}{\rm d}x\left\{\frac{{\rm i}\pi^{2}}{m_{5}^{2}-m_{12x}^{2}}\ln\frac{m_{12x}^{2}}{m_{5}^{2}}-\left[\left(q_{1}-\bar{x}k\right)^{2}+m_{5}^{2}-m_{12x}^{2}\right]\pi\bar{J}_{1}^{(2)}\left(2;1,1\right)\right\}.$ (73)
Note that at $\varepsilon=0$ the imaginary part of ${\rm i}^{-1}J^{(2)}(2-2\varepsilon;1,1)$ has a simple structure:
${\rm Im}\left[{\rm i}^{-1}J^{(2)}\left(2;1,1\right)\right]_{s}=\frac{2\pi^{2}}{\sqrt{\left(s-m_{12x}^{2}-m_{5}^{2}\right)^{2}-4m_{12x}^{2}m_{5}^{2}}},$
At this point, we are ready to complete integration over second loop momentum $q_{1}$:
@@
$+\frac{1}{\pi^{2}}\int\limits_{(m_{5}+m_{12x})^{2}}^{\infty}{\rm d}s\;\frac{{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2;1,1\right)\right]_{s}}{s}\left[\left(s+m_{5}^{2}-m_{12x}^{2}\right)C_{0}+B_{0}\right]\Biggr\},$ (74)
@@
$U^{(2)}=\frac{1}{2\pi^{2}}\int\limits_{0}^{1}{\rm d}x\int\limits_{(m_{5}+m_{12x})^{2}}^{\infty}\frac{{\rm d}s}{s}\left(s+m_{5}^{2}-m_{12x}^{2}\right)\;{\rm{Im}}\left[{\rm i}^{-1}J^{(2)}\left(2;1,1\right)\right]_{s}C_{0}.$ (75)